# MotionCharacter: Identity-Preserving and Motion Controllable Human Video Generation

### v2

@@ Abstract
Recent advancements in personalized Text-to-Video (T2V) generation highlight the importance of integrating character-specific identities and actions. However, previous T2V models struggle with identity consistency and controllable motion dynamics, mainly due to limited fine-grained facial and action-based textual prompts, and datasets that overlook key human attributes and actions. To address these challenges, we propose MotionCharacter, an efficient and high-fidelity human video generation framework designed for identity preservation and fine-grained motion control. We introduce an ID-preserving module to maintain identity fidelity while allowing flexible attribute modifications, and further integrate ID-consistency and region-aware loss mechanisms, significantly enhancing identity consistency and detail fidelity. Additionally, our approach incorporates a motion control module that prioritizes action-related text while maintaining subject consistency, along with a dataset, Human-Motion, which utilizes large language models to generate detailed motion descriptions. For simplify user control during inference, we parameterize motion intensity through a single coefficient, allowing for easy adjustments. Extensive experiments highlight the effectiveness of MotionCharacter, demonstrating significant improvements in ID-preserving, high-quality video generation.
Figure 1: Given a single reference facial image, MotionCharacter can generate identity-consistent video outputs across text prompts, action phrases, and motion intensities. The upper section demonstrates its capability to accurately follow specific action phrases, while the lower section highlights its fine-grained motion control achieved by varying user-defined motion intensities.
1 Introduction
@@
High-quality, personalized, and controllable human video generation has gained significant traction, with applications spanning social media, virtual avatars, and personalized content creation. Recent advancements in text-driven video generation models [15, 9, 14, 28, 3, 32, 35, 44, 4, 5, 2] have driven substantial progress in this field, yet several challenges remain, particularly in maintaining identity consistency and achieving fine-grained control over motion instructions. In response, recent approaches [19, 36, 23, 38, 11] in subject-driven Text-to-Video (T2V) generation have explored ways to address the challenge of producing high-quality videos that accurately depict specific individuals while following motion instructions consistently.
However, most of these approaches rely on separate training for each identity (ID), limiting their scalability and flexibility in practical applications. Some works such as ID-Animator [11] have advanced the field by enabling identity-specific video generation using any reference facial image without requiring additional fine-tuning, but it still lacks the ability to finely control the intensity of motion, constraining its responsiveness to nuanced motion prompts. This limitation underscores the need for methods that can offer both identity preservation and precise motion control in T2V generation. Besides, the typical prompts provided by users to T2V models include descriptions of the entire video content, encompassing scene details and human motion. However, text-based prompts alone are often insufficient to capture fine-grained motion dynamics accurately. For example, phrases like “open mouth” or “eyes closed” omit key details such as movement speed or intensity, which are essential for nuanced motion dynamics. While a more intuitive approach might be to use motion-only prompts, current T2V models exhibit limited sensitivity to such concise motion instructions, often failing to reflect the intended subtleties in generated videos. In addition, the existing facial text-video datasets such as CelebV-Text [43] primarily focus on emotional changes, neglecting essential human attributes and actions, rendering them inadequate for identity-preserving video generation tasks.
In this paper, we propose MotionCharacter, a human video generation framework specifically designed for identity preservation and fine-grained motion control. To achieve high-fidelity and identity consistency for human video generation, we introduce an ID-preserving module and employ a combination of face embedding and CLIP [25] embedding, allowing the model to retain high identity fidelity while also being flexible enough to accommodate dynamic modifications of attributes such as expressions or actions based on user prompts. Additionally, we integrate a composite loss function that incorporates both ID-consistency loss and region-aware loss components that direct the model’s attention to critical facial regions, addressing common issues like distortion or blurriness in features such as lips and teeth, and enhance identity consistency and detail fidelity in personalized T2V generation.
To enhance the model’s responsiveness to motion instructions, we introduce a specialized motion control module that prioritizes action-related text cues while preserving identity consistency throughout the video sequence. Complementing this module, we present a new dataset, Human-Motion, which employs large language models (LLMs) to produce comprehensive and nuanced motion descriptions tailored for identity-preserving video synthesis. Furthermore, to facilitate user control over motion dynamics, we introduce a parameterized motion intensity coefficient, allowing users to easily adjust the scale of movement during inference. Together, these advancements improve the model’s ability to accurately follow motion instructions and generate realistic, personalized content. Through extensive experiments, we present qualitative, quantitative, and user study results that validate the effectiveness of our method in terms of identity consistency and adherence to motion instructions. In summary, our contributions are as follows:
  • We propose a framework, named MotionCharacter, designed to enhance identity consistency and fine-grained motion controllability in human video generation.
  • We introduce a motion control module that prioritizes action-related text while maintaining subject consistency, which enables more precise control over motion dynamics and improves the generation of high-quality human videos based on textual descriptions.
  • We propose region-aware loss to improve attention to critical facial regions, ensuring high-quality and identity-preserving video generation while maintaining accurate motion dynamics.
2 Related Work
@@
Text-to-Video Diffusion Model. Recent advancements in diffusion models [13, 29, 26] have positioned them as a prominent method in generative modeling, especially in text-to-video (T2V) generation. The Video Diffusion Model [15] was among the first to utilize a space-time-factored U-Net architecture for unconditional video generation, effectively modeling video distributions in pixel space. Building on this, AnimateDiff [9] advanced the field by incorporating a motion module into the Stable Diffusion framework, enhancing its ability to generate videos from textual prompts. Subsequent developments included Imagen Video [14] and Make-a-Video [28], which introduced sequential models for T2V generation, focusing on pixel-space representations. In response to the challenges associated with high-dimensional video data, latent video diffusion model [3] proposed leveraging latent diffusion models to operate within the latent space of an auto-encoder. This latent approach has gained traction, leading to a proliferation of methods such as ModelScope [32], LAVIE[35], MagicVideo [44], VideoCrafter [4, 5], each contributing to the ongoing evolution of T2V models and their applications.
Identity-Preserving Image Generation. The task of identity-preserving image generation aims to synthesize visual content that maintains the unique characteristics of a specific individual while allowing for variations in pose, expression, and other attributes. Most techniques focus on facial images, using methods like texture-based approaches and latent space manipulation to preserve identity. While fine-tuning methods such as Low-Rank Adaptation [16], Textual Inversion [7], and DreamBooth [27] can customize models for ID-specific images, they often require dedicated training for each identity. Recent approaches utilize embeddings as conditional inputs to guide the generation process, enabling explicit identity control. For example, IP-Adapter [42] integrates an adapter module with identity embeddings into a pre-trained model for efficient identity preservation. PhotoMaker [20] stacks multiple images to mitigate identity-irrelevant features. InstantID [33] allows real-time, identity-specific generation with a lightweight architecture, reducing the need for case-specific training. PuLID [10] employs a multi-stage process to refine identity embeddings, enhancing fidelity and visual quality.
Subject-Driven Text-to-Video Generation. Subject-driven text-to-video generation aims to incorporate specific characters or subjects into synthesized videos while allowing for text-based control over actions, styles, and sequences. Previous methods such as VideoBooth [19], DreamVideo [36], MagicMe [23], and CustomCrafter [38] have explored learning-based frameworks to combine visual identity with motion dynamics. However, these approaches often require separate training for each individual, which can limit their scalability and flexibility. In contrast, the recent development of ID-Animator [11] has demonstrated the ability to achieve zero-shot, training-free capabilities, allowing for identity incorporation without extensive retraining. Nonetheless, this method still lacks fine-grained control over motion intensity. Specifically, the inability to adjust the strength of movements can hinder the model’s effectiveness in generating realistic and dynamic animations that align with user specifications. In response to these limitations, our approach, MotionCharacter, enables simultaneous control of both appearance and motion without necessitating retraining during inference, thereby enhancing the usability and efficiency of subject-driven video generation.
@@
Figure 2: Framework overview. Our proposed framework comprises three core components: the ID-Preserving Module, the Motion Control Module, and a composite loss function. The loss function incorporates a Region-Aware Loss to ensure high motion fidelity and an ID-Consistency Loss to maintain alignment with the reference ID image. During training, motion intensity $\mathcal{M}$ is derived from optical flow. At inference, human animations are generated based on user-defined motion intensity $\mathcal{M}$ and specified action phrases, enabling fine-grained and controllable video synthesis.
3 Methodology
3.1 Problem Formulation
@@
Personalized human video generation aims to create vivid clips consistent in character identity and motion based on a reference image and text prompt. To achieve this goal, we propose a novel model named MotionCharacter which accurately reflects identity information, captures action-based motion, and maintains smooth visual transitions. Formally, given a reference ID image $\mathcal{I}$, a text prompt $\mathcal{P}$, an action phrase $\mathcal{A}$, and a motion intensity $\mathcal{M}$, the model $\mathcal{F}$ is designed to produce video $\mathcal{V}$ by:
@@
Technically, we elaborately design the structure from two aspects, ID-Preserving Optimization and Motion Control Enhancement. Besides, we propose a new Human Motion dataset which is specially curated and annotated for training high-fidelity human video generation.
3.2 ID-Preserving Optimization
ID Content Insertion. Since the adopted pretrained text-to-video (T2V) diffusion model [9] lacks identity-preserving capabilities, we first intend to introduce an ID-Preserving Adapter into the backbone to emphasize identity-specific regions and reduce irrelevant background interference. As illustrated in Fig. 2, the ID-Preserving Adapter extracts the identity embedding $C_{id}$ from the reference image $\mathcal{I}$ and injects the identity embedding $C_{id}$ into the diffusion model through cross-attention.
Specifically, the face region is first isolated from the reference image $\mathcal{I}$ to filter the interference of the background region. Then the face region image is processed in parallel to a pre-trained CLIP image encoder [25] and a face recognition model ArcFace [6] to obtain the broad contextual identity embeddings $E_{clip}$ and the fine-grained identity embeddings $E_{arc}$, respectively. To effectively combine global context with fine-grained identity details, we employ cross-attention to fuse the CLIP and ArcFace embeddings:
@@ 3.2 ID-Preserving Optimization
where $W_{q}^{\prime}$, $W_{k}^{\prime}$, and $W_{v}^{\prime}$ are learnable parameters, with $E_{arc}$ as the query and the combined embedding $E=E_{clip}+E_{arc}$ as the key and value. Following cross-attention, a projection layer Proj is applied to align the dimension with the text embedding, thereby generating the final identity embedding $C_{id}$ for the reference image $\mathcal{I}$.
Inspired by recent work on image prompt adapters [42, 33], the identity embedding $C_{id}$ in MotionCharacter is regarded as an image prompt embedding and is used alongside text prompt embeddings to provide guidance for the diffusion model. This procedure can be expressed as:
@@
ID-Consistency Loss. Current stable diffusion models are usually trained with the MSE loss over each pixel, which is insufficient to ensure identity preservation between the reference image and the generated videos. To address this issue, we introduce an ID-Consistency loss during training phase to maintain the identity information.
Specifically, at a specific diffusion step $\hat{t}$, the diffusion model can estimate the noise-free latent ${\hat{z}}_{0}$ from a noisy latent $z_{\hat{t}}$ by DDIM reversion process. Then, the estimated $\hat{z}_{0}$ is passed to a VAE decoder to reconstruct the frame, denoted as $X^{f}$. Therefore, the ID-Consistency loss $\mathcal{L}_{id}$ across the sequence of $N$ frames can be calculated by:
@@
where $\cdot$ denotes the dot product, $\phi$ denotes the face recognition backbone [6], $\phi(X^{f}_{i})$ and $\phi(I)$ represent the face embedding of each generated frame $i$ and the reference identity image $I$, respectively.
3.3 Motion Control Enhancement
Although text prompt embedding shows control capacity to some degree, it is still insufficient to capture fine-grained motion dynamics. To address this challenge, we propose a spatial-aware motion control module with motion intensity to enhance the controllability. Besides, a region-aware loss is employed to enhance spatial coherence and realism in dynamic regions such as the face.
Motion Control Module. We regard the control capacity of the model as lying in two aspects: one is the faithfulness of the motion description, and the other is the magnitude of motion intensity. To achieve this goal, we introduce extra action phrase and motion intensity as the conditions in the proposed model.
We first extract the action phrase $\mathcal{A}$ from the original text prompt $\mathcal{P}$, for example, “talking” from “a man is talking”. We resort to MiniGPT [45] to implement this extraction process automatically. Then the action phrase is fed to CLIP text encoder [25] to obtain the action embedding $E_{M}$ which captures the semantic intent of the motion.
Considering the magnitude of motion intensity is hard to define directly, we employ an optical flow estimator to extract the optical flow magnitude of the video as the motion intensity. Specifically, given a video clip $\mathcal{V}^{in}=\{v^{in}_{i}\}_{i=1}^{N}$, where $N$ is the number of frames, we first extract the optical flow of each pixel between two adjacent frames by:
@@ 3.3 Motion Control Enhancement
where $(x,y)$ denotes the position of each pixel, and $\Theta$ is an optical flow estimation model. We use RAFT [30] as $\Theta$ for efficient and accurate optical flow estimation. Then the mean optical flow value $\tau_{i}$ can be calculated by simply averaging $f_{i,(x,y)}$. Afterward, we take $\tau_{i}$ as the threshold to produce binary mask $M_{i,(x,y)}$. Specifically, when the magnitude of the optical flow exceeds $\tau_{i}$, set the corresponding position in $M_{i,(x,y)}$ to 1; otherwise set it to 0. Consequently, the mean foreground optical flow value $f_{i,fg}$ can be easily obtained by:
@@
where $f_{i,fg}(x,y)$ is the foreground optical flow at each pixel $(x,y)$. $S$ denotes the number of the foreground pixels. The motion intensity $\mathcal{M}$ of the video is defined as follows:
$\mathcal{M}=\frac{1}{N-1}\sum^{N-1}_{i=0}f_{i,fg}.$ (7)
Subsequently, motion intensity $\mathcal{M}$ is projected through a multi-layer perceptron (MLP) to generate a motion embedding $E_{M}$ aligned with the dimensionality of the action embedding $E_{A}$.
As illustrated in Fig. 2, two parallel cross attention modules (Cross Attn and Motion Attn) are adopted in the motion control module to insert the action embedding $E_{A}$ and motion embedding $E_{M}$. The process is formally represented as follows:
@@
where $Q^{\prime}=Z^{\prime}W^{\prime}_{q}$ is relevant with the output of ID-Preserving Module. $K^{a}$, $V^{a}$ and $K^{m}$, $V^{m}$ are the key-value pairs derived from the action embedding $E_{A}$ and the motion embedding $E_{M}$, respectively. The parameter $\alpha$ balances the influence of motion intensity within the combined attention output $Z^{\prime\prime}$.
[…]
IPA-PlusFace[42] 0.797 0.985 0.325 0.587 0.218 0.480
IPA-FaceID-Portrait[42] 0.849 0.984 0.191 0.545 0.223 0.531
IPA-FaceID-PlusV2[42] 0.813 0.987 0.085 0.575 0.217 0.617
ID-Animator[11] 0.857 0.979 0.433 0.607 0.204 0.546
Ours 0.869 0.998 0.449 0.633 0.227 0.609
Table 1: Comparison of different methods across multiple metrics. Higher values indicate better performance, with the best scores in bold and the second best in underline. It is important to note that all methods were configured with an empty action phrase and a motion intensity setting of 20 for a more dynamic effect.
Region-Aware Loss. The fluency of the generated video heavily relies on the spatial coherence and realism of dynamic regions, e.g. the face areas. To achieve this goal, we apply a region-aware loss to force the model to focus more on the high-motion regions. Specifically, we normalize the foreground optical flow $f_{i,fg}(x,y)$ defined in Eq. (6) and calculate the optical flow mask $M_{i,\text{norm}}$:
$M_{i,\text{norm}}=clip\left(f_{i,fg}(x,y)/255+0.5,\,1.0,\,1.5\right),$ (9)
where $clip(\cdot,\,a,\,b)$ restricts the values into $[a,b]$. The high-motion areas will be assigned a greater value than the low-motion regions. Then the region-aware loss $\mathcal{L}_{R}$ across all $N$ frames can be compactly defined as:

### v3

# MotionCharacter: Fine-Grained Motion Controllable Human Video Generation

@@ Abstract
Recent advancements in personalized Text-to-Video (T2V) generation have made significant strides in synthesizing character-specific content. However, these methods face a critical limitation: the inability to perform fine-grained control over motion intensity. This limitation stems from an inherent entanglement of action semantics and their corresponding magnitudes within coarse textual descriptions, hindering the generation of nuanced human videos and limiting their applicability in scenarios demanding high precision, such as animating virtual avatars or synthesizing subtle micro-expressions. Furthermore, existing approaches often struggle to preserve high identity fidelity when other attributes are modified. To address these challenges, we introduce MotionCharacter, a framework for high-fidelity human video generation with precise motion control. At its core, MotionCharacter explicitly decouples motion into two independently controllable components: action type and motion intensity. This is achieved through two key technical contributions: (1) a Motion Control Module that leverages textual phrases to specify the action type and a quantifiable metric derived from optical flow to modulate its intensity, guided by a region-aware loss that localizes motion to relevant subject areas; and (2) an ID Content Insertion Module coupled with an ID-Consistency loss to ensure robust identity preservation during dynamic motions. To facilitate training for such fine-grained control, we also curate Human-Motion, a new large-scale dataset with detailed annotations for both motion and facial features. Extensive experiments demonstrate that MotionCharacter achieves substantial improvements over existing methods. Our framework excels in generating videos that are not only identity-consistent but also precisely adhere to specified motion types and intensities.
Project Pagehttps://motioncharacter.github.io/
Introduction
Figure 1: Comparison of subject-driven T2V methods and our proposed MotionCharacter framework. Existing approaches specify coarse actions (e.g., “open mouth”) but struggle to capture nuanced motion magnitude (e.g., “slightly” vs. “widely”). In contrast, MotionCharacter decouples action type and motion intensity, enabling fine-grained, continuous control over human motion while preserving subject fidelity.
@@ Introduction
High-quality, personalized, and controllable human video generation has gained significant traction, with applications spanning social media, film production, virtual avatars, and personalized content creation. Recent pioneering advancements in text-driven video generation models (ho2022video; guo2023animatediff; ho2022imagen; wang2023modelscope; wang2023lavie; zhou2022magicvideo; chen2024videocrafter2) have driven substantial progress in this field. In particular, subject-driven Text-to-Video (T2V) generation models (jiang2024videobooth; wei2024dreamvideo; ma2024magic; wu2024customcrafter; he2024id) have made strides in producing high-quality videos that faithfully depict specific individuals.
However, existing subject-driven T2V models still face a significant limitation: a lack of fine-grained control over motion. This deficiency curtails their applicability in scenarios demanding high precision, such as animating digital humans with precise expressions, synthesizing subtle micro-expressions for psychological analysis, or creating realistic virtual avatars that respond dynamically to user input.
The core of this problem lies in the entanglement of action semantics and motion intensity within textual prompts. As shown in Fig. 1(a), current subject-driven T2V models (he2024id; guo2023animatediff) allow users to specify human actions using action phrases like “open mouth”. However, these phrases provide only a coarse description and fall short in precisely controlling motion intensity. Even when employing more nuanced phrases (e.g., “open mouth slightly”, “open mouth widely”), the generated results often lack the intended subtleties because language captures actions in a discrete manner, whereas motion intensity is inherently continuous. This forces the model to guess the intended magnitude, leading to unpredictable and uncontrollable results.
To break this entanglement and enable fine-grained controllability, we introduce MotionCharacter, a framework that explicitly decouples motion into two components: action type and motion intensity. To achieve this, we first propose a Motion Control Module. This module uses simple action phrases (e.g., “smile”) to define the type of motion, while a separate, continuous signal derived from optical flow modulates its intensity. This allows a user to specify an action with text and then fine-tune its magnitude precisely, for instance, via a simple slider. To further enhance motion dynamics and the quality of facial transitions, this module is enhanced by a region-aware loss that directs the model’s attention to critical facial regions like the lips and eyes, preventing distortion during motion.
Moreover, as motions become more dynamic and expressive, maintaining the subject’s identity becomes a significant challenge. To address this, we introduce an ID Preservation Module that injects detailed facial features into the generation process. This is reinforced by a powerful ID consistency loss to ensure that the character’s core identity remains stable and detailed, even during large and complex movements.
Finally, while several human video datasets (zhu2022celebv; yu2023celebv) exist, they predominantly focus on capturing basic emotional expressions or broad action categories, lacking the granular annotations needed for fine-grained motion control. Also, these datasets are limited in scale and contain unfiltered multi-subject and low-quality scenes. This limitation creates a significant bottleneck in training models capable of nuanced motion generation. To address this gap, we introduce Human-Motion, a large-scale dataset of 106,292 video clips. We use Large Multimodal Models (LMMs) and optical flow estimation to generate detailed annotations for motion type and intensity, aiming to facilitate research on fine-grained human video generation.
Through extensive experiments, we present qualitative, quantitative, and user study results that validate the performance of our method in terms of identity consistency and adherence to motion instructions. In summary, our contributions are as follows:
  1. 1.
    We propose MotionCharacter, a framework for personalized human video generation that decouples motion control into action type and intensity, enabling fine-grained and continuous motion adjustments.
  2. 2.
    We introduce a Motion Control Module that leverages text for action semantics and a continuous signal for intensity, enhanced by a region-aware loss to ensure high-quality dynamics. We also propose an ID Preservation Module with a dedicated ID-Consistency loss to maintain robust identity fidelity.
  3. 3.
    We curate Human-Motion, a large-scale and high-quality dataset of 106,292 video clips with detailed annotations for motion and identity, created specifically to facilitate research on controllable human video generation.
Related Work
@@
Text-to-Video Diffusion Model. Recent advancements in diffusion models (ho2020denoising; song2020score; rombach2022high) have established them as a leading approach in Text-to-Video generation. The Video Diffusion Model (ho2022video) pioneered a space-time-factored U-Net for unconditional video generation in pixel space. Building on this, AnimateDiff (guo2023animatediff) integrated a motion module into the Stable Diffusion framework to generate videos from text prompts. Later works, such as Imagen Video (ho2022imagen) and Make-a-Video (singer2022make), adopted sequential models for pixel-space T2V generation. To overcome challenges with high-dimensional video data, latent diffusion models (blattmann2023align) operate in the latent space of an auto-encoder, inspiring further developments including ModelScope (wang2023modelscope), LAVIE (wang2023lavie), MagicVideo (zhou2022magicvideo), and VideoCrafter (chen2023videocrafter1; chen2024videocrafter2).
Subject-Driven Image and Video Generation. Subject-driven generation aims to synthesize content while preserving specific subject characteristics. In the image domain, recent works like IP-Adapter (ye2023ip), PhotoMaker (li2024photomaker), and InstantID (wang2024instantid) have achieved efficient identity preservation through embedding-based approaches, eliminating the need for subject-specific training required by methods like DreamBooth (ruiz2023dreambooth). For video generation, works such as VideoBooth (jiang2024videobooth), DreamVideo (wei2024dreamvideo), and CustomCrafter (wu2024customcrafter) have explored learning-based frameworks to combine visual identity with motion dynamics. While ID-Animator (he2024id) demonstrates zero-shot capabilities, it lacks fine-grained control over motion intensity. Our MotionCharacter addresses these limitations by enabling control of both appearance and motion without requiring retraining during inference.
@@
Figure 2: Framework overview. Our proposed framework comprises three core components: the ID Content Insertion Module, the Motion Control Module, and a composite loss function. The loss function incorporates a Region-Aware Loss to ensure high motion fidelity and an ID-Consistency Loss to maintain alignment with the reference ID image. During training, motion intensity $\mathcal{M}$ is derived from optical flow. At inference, human animations are generated based on user-defined motion intensity $\mathcal{M}$ and specified action phrases, enabling fine-grained and controllable video synthesis.
Methodology
@@
Overall Pipeline. Personalized human video generation aims to create vivid clips consistent in character identity and motion based on a reference image and text prompt. To achieve this goal, we propose a framework named MotionCharacter which accurately reflects identity information, captures fine-grained motion and maintains smooth visual transitions. Our framework is shown in Fig. 2. Formally, given a reference ID image $\mathcal{I}$, a text prompt $\mathcal{P}$, an action phrase $\mathcal{A}$, and a motion intensity $\mathcal{M}$, the model $\mathcal{F}$ is designed to produce video $\mathcal{V}$ by:
@@
Our methodology is built upon the central principle of disentanglement, designed to establish an orthogonal control space for human video generation. We recognize that to robustly control motion, one must first guarantee the stability of identity, as intense dynamics often corrupt subject-specific features. Therefore, our framework is architected around two synergistic pillars that address this challenge sequentially. First, our ID-Preserving Optimization strategy establishes a stable identity foundation. Second, building upon this foundation, our Motion Control Enhancement module operates to independently manipulate action type and intensity. Furthermore, to enable robust training for such a disentangled framework, we introduce the Human-Motion dataset, specifically curated with dual-track annotations for motion semantics and intensity.
ID-Preserving Optimization
ID Content Insertion. Since the adopted pretrained Text-to-Video (T2V) diffusion model (guo2023animatediff) lacks identity-preserving capabilities, we first intend to introduce an ID Content Insertion Module into the backbone to emphasize identity-specific regions and reduce irrelevant background interference. As illustrated in Fig. 2, the module extracts the identity embedding $C_{id}$ from the reference image $\mathcal{I}$ and injects the identity embedding $C_{id}$ into the diffusion model through cross-attention.
Specifically, the face region is first isolated from the reference image $\mathcal{I}$ to filter the interference of the background region. Then the face region image is processed in parallel to a pre-trained CLIP image encoder (radford2021learning) and a face recognition model ArcFace (deng2019arcface) to obtain the broad contextual identity embeddings $E_{clip}$ and the fine-grained identity embeddings $E_{arc}$, respectively. To effectively combine global context with fine-grained identity details, we employ cross-attention to fuse the CLIP and ArcFace embeddings:
@@
where $W_{q}^{\prime}$, $W_{k}^{\prime}$, and $W_{v}^{\prime}$ are learnable parameters, with $E_{arc}$ as the query and the combined embedding $E=E_{clip}+E_{arc}$ as the key and value. This fusion allows the detailed ArcFace features to selectively attend to the most relevant contextual information from CLIP embeddings. Following cross-attention, a projection layer Proj is applied to align the dimension with the text embedding, thereby generating the final identity embedding $C_{id}$ for the reference image $\mathcal{I}$.
Inspired by recent work on image prompt adapters (ye2023ip; wang2024instantid), the identity embedding $C_{id}$ in MotionCharacter is regarded as an image prompt embedding and is used alongside text prompt embeddings to provide guidance for the diffusion model. This procedure can be expressed as:
@@
ID-Consistency Loss. To enforce identity preservation at a semantic level, we introduce an ID-Consistency loss that complements the standard pixel-wise MSE objective. The MSE loss, while crucial for visual fidelity, is fundamentally agnostic to high-level concepts like identity. Our proposed loss addresses this by penalizing deviations directly within a learned identity feature space. Specifically, it minimizes the feature-space distance between the reference identity and the generated frames, ensuring that the model maintains the subject’s core characteristics even during complex motion.
In practice, at a specific diffusion step $\hat{t}$, the diffusion model can estimate the noise-free latent ${\hat{z}}_{0}$ from a noisy latent $z_{\hat{t}}$ via the DDIM reversion process. Then, the estimated $\hat{z}_{0}$ is passed to a VAE decoder to reconstruct the frame, denoted as $X^{f}$. Therefore, the ID-Consistency loss $\mathcal{L}_{id}$ across the sequence of $N$ frames can be calculated by:
@@
where $\cdot$ denotes the dot product, $\phi$ denotes the pre-trained face recognition backbone (deng2019arcface), $\phi(X^{f}_{i})$ and $\phi(I)$ represent the normalized face embedding of each generated frame $i$ and the reference identity image $I$, respectively. $N$ is the total number of frames.
Motion Control Enhancement
In subject-driven T2V models, achieving precise motion control remains a significant challenge. While previous works (he2024id) have successfully generated identity-specific videos, they often fail to capture fine-grained motion dynamics, resulting in limited responsiveness to subtle motion cues. To address this limitation, we propose a spatial-aware motion control module that explicitly incorporates motion intensity information to enhance controllability. Moreover, a region-aware loss is employed to enhance spatial coherence and realism in dynamic regions such as face.
We regard the control capacity of the model as lying in two aspects: one is the faithfulness of the motion description, and the other is the magnitude of motion intensity. To achieve this goal, we introduce extra action phrase and motion intensity as the conditions in the proposed model. We first use LMM (zhu2023minigpt) to automatically generate overall descriptions and action phrases, enriching the dataset with motion-related information and serving as the primary text description $\mathcal{P}$ and action phrase $\mathcal{A}$, as defined in Eq. 1. Then the action phrase is fed to CLIP text encoder (radford2021learning) to obtain the action embedding $E_{A}$ which captures the semantic intent of the motion.
Motion Intensity Estimation. The magnitude of motion intensity is challenging to define directly, especially considering the diversity of movements and their varying characteristics. To address this, we employ optical flow estimation to capture motion intensity. Optical flow captures pixel-wise motion between adjacent frames, offering a fine-grained measure of movement that directly reflects the motion of objects in the video. Specifically, given a video clip $\mathcal{V}^{in}=\{v^{in}_{i}\}_{i=1}^{N}$, where $N$ is the number of frames, we first extract the optical flow of each pixel between two adjacent frames by:
@@
where $(x,y)$ denotes the position of each pixel, and $\Theta$ is an optical flow estimation model. We use RAFT (teed2020raft) as $\Theta$ for efficient and accurate optical flow estimation. Then the mean optical flow value $\tau_{i}$ can be calculated by simply averaging $f_{i,(x,y)}$. Afterward, we take $\tau_{i}$ as the threshold to produce binary mask $M_{i,(x,y)}$. Specifically, when the magnitude of the optical flow exceeds $\tau_{i}$, set the corresponding position in $M_{i,(x,y)}$ to 1; otherwise set it to 0. Consequently, the mean foreground optical flow value $f_{i,fg}$ can be obtained by:
@@
where $f_{i,fg}(x,y)$ is the foreground optical flow at each pixel $(x,y)$. $S$ denotes the number of the foreground pixels.
The motion intensity $\mathcal{M}$ of the video is then defined as the average foreground optical flow value across all frames:
$\mathcal{M}=\frac{1}{N-1}\sum^{N-1}_{i=1}f_{i,fg},$ (7)
where $\mathcal{M}$ is a scalar representing the motion intensity and $N$ is the number of frames. Subsequently, the motion intensity $\mathcal{M}$ is passed through a motion intensity estimator, utilizing a multi-layer perceptron (MLP) to generate a motion intensity embedding $E_{M}$ aligned with the dimensionality of the action embedding $E_{A}$.
Motion Condition Injection. As illustrated in Fig. 2, two parallel cross attention modules are adopted in the motion control module to insert the action embedding $E_{A}$ and motion intensity embedding $E_{M}$. The process is formally represented as follows:
@@
where $Q^{\prime}=Z^{\prime}W^{\prime}_{q}$ is relevant with the output of ID Content Insertion Module. $K^{a}$, $V^{a}$ and $K^{m}$, $V^{m}$ are the key-value pairs derived from the action embedding $E_{A}$ and the motion intensity embedding $E_{M}$, respectively. The parameter $\alpha$ balances the influence of motion intensity within the combined attention output $Z^{\prime\prime}$ and is set to 1 by default.
Although both our approach and Stable Video Diffusion (SVD) (blattmann2023stable) leverage optical flow to characterize motion, a key distinction lies in our explicit decoupling of motion cues. SVD uses motion bucket parameter, which is integrated via additional time embeddings, to coarsely control overall motion magnitude. This results in a global and non-localized representation that may obscure subtle variations (e.g., faces). In contrast, our dual-branch motion condition injection strategy separates semantic guidance (what motion is desired) from quantitative strength (how intense it should be). The use of parallel cross-attention to fuse these decoupled cues is a key mechanism that enables predictable and fine-grained control, allowing our model to capture subtle motion nuances with high fidelity.
Region-Aware Loss. The fluency of the generated video heavily relies on the spatial coherence and realism of dynamic regions, e.g., the face areas. To achieve this goal, we apply a region-aware loss to force the model to focus more on the high-motion regions. Specifically, we normalize the foreground optical flow $f_{i,fg}(x,y)$ defined in Eq. (6) and calculate the optical flow mask $M_{i,\text{norm}}$:
$M_{i,\text{norm}}=\text{clip}\left(\frac{f_{i,fg}(x,y)}{255}+\delta,\,1.0,\,1.0+\delta\right),$ (9)
where $\text{clip}(\cdot,\,a,\,b)$ restricts the values to $[a,b]$ and $\delta$ is a scalar offset that modulates the spatial weighting of the optical flow mask. This mask assigns higher weights to regions with significant motion, ensuring that both the primary subject (e.g., faces) and high-dynamic background regions, especially when the text prompt describes a moving background, receive adequate attention. Note that although the optical flow mask assigns non-zero weights to some background regions, it ensures that regions with significant motion from both the subject and background are appropriately emphasized in the loss computation. In contrast to segmentation, depth maps and other similar static scene cues that may overlook such background dynamics, the optical flow mask retains comprehensive motion information across the scene. Then the region-aware loss $\mathcal{L}_{R}$ across all $N$ frames can be compactly defined as: